I need your help
Solution A is 30% acid and solution B is 60% acid. How many gallons of each solution must be used to create a 60 gallons miture that is 50% acid?
Thanks

Answer by nerdybill(7384)

Solution A is 30% acid and solution B is 60% acid. How many gallons of each solution must be used to create a 60 gallons miture that is 50% acid?
.
Let x = amount (gal) of 30%
then
60-x = amount of 60%
.
.30x + .60(60-x) = .50(60)
.30x + 36-.60x = 30
36-.30x = 30
-.30x = -6
x = 20 gals of 30% acid
.
60% acid:
60-x = 60-20 = 40 gals